Each business day, on average, a company writes checks totaling $13,400 to pay its suppliers. The usual clearing time for the checks is four days. Meanwhile, the company is receiving payments from its customers each day, in the form of checks, totaling $24,400. The cash from the payments is available to the firm after two days.

a. Calculate the company’s disbursement float, collection float, and net float. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

b. If the collected funds were available in one day instead of two, what would be the company's disbursement float, collection float, and net float? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)
Answer #1

a. Disbursement float = 4 days * 13400 = 53600

Collection float = 2 days * ( -24400) = - 48800

Net float = 53600 - 48800 = 4800

b. Disbursement float = 4 days * 13400 = 53600

Collection float = 1 days * ( -24400) = - 24400

Net float = 53600 - 24400 = 29200